# Analytics, have some humility: a statistical view of fourth-down decision making

### v5

@@ Appendix C Tuning the number of bootstrapped datasets
We use ${\mathsf{boot\%}}$ to measure uncertainty in the estimated optimal decision, but we cannot necessarily rely on a non-technical football coach to fully process how ${\mathsf{boot\%}}$ maps to the strength of a recommendation. To facilitate easier communication to a coach, we bin decision recommendations into three buckets: confident (${\mathsf{boot\%}}\in[90\%,100\%]$), lean (${\mathsf{boot\%}}\in[75\%,90\%)$), and uncertain (${\mathsf{boot\%}}\in[0\%,75\%]$). Since we ultimately bucket decision recommendations into these three bins (see Section 4.2), we want the number of bootstrapped models $B$ to be large enough to stably categorize decisions into one of these three bins. We also want $B$ to be as small as possible in order to quickly evaluate a fourth-down recommendation during a football game. Hence, we conduct a stability analysis to choose $B$.
@@
$\begin{cases}p_{i}^{\text{confident, }(B)}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left({\mathsf{boot\%}}_{im}^{(B)}\in[90\%,100\%]\right),\\ p_{i}^{\text{lean, }(B)}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left({\mathsf{boot\%}}_{im}^{(B)}\in[75\%,90\%)\right),\\ p_{i}^{\text{uncertain, }(B)}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left({\mathsf{boot\%}}_{im}^{(B)}\in[0\%,75\%)\right),\\ p_{i}^{(B)}=\max\{p_{i}^{\text{confident, }(B)},\ p_{i}^{\text{lean, }(B)},\ p_{i}^{\text{uncertain, }(B)}\}.\end{cases}$ (27)
@@ Appendix D Baseline coaches’ decision model
To compare our decision-making procedure to the decisions that actual football coaches tend to make, we model the probability that a coach chooses a decision in $\{{\mathsf{Go}},{\mathsf{FG}},{\mathsf{Punt}}\}$ as a function of game-state. We use ${\mathsf{XGBoost}}$ to fit these coach probabilities. ${\mathsf{XGBoost}}$ works well here because we have $94,786$ fourth-down plays in our full dataset of plays since 1999, and each play is a reasonably approximate independent observation of a coach’s decision. In particular, we fit these coach probabilities as a function of yardline, yards to go, game seconds remaining, score differential, point spread, and era (1999-2001, 2002-2005, 2006-2013, 2014-2017, and 2018-present). In Figure 12 we visualize these coach decision models, and the results make intuitive sense. For the most part, coaches punt deep in their own territory and kick field goals near the opponent’s endzone, except for with one and sometimes two yards to go. Also, at the end of the game, coaches’ decision-making changes depending on the number of points they need to score to win the game.
Figure 12: Visualizing our model of the typical coach’s fourth-down decision as a function of yardline and yards to go, for various values of time remaining and score differential. Green, yellow, and red indicate that ${\mathsf{Go}}$, ${\mathsf{FG}}$, and ${\mathsf{Punt}}$ is the most likely decision, respectively. The color intensity reflects the likelihood that a coach makes that decision.
@@ Appendix E Additional example plays
Example play 3. In Figure 14 we visualize the decision procedure for a fourth-down play in which the Bears had the ball against the Jets in Week 12 of 2022. ${\mathsf{FG}}$ provides an edge over ${\mathsf{Go}}$ according to the ${\mathsf{WP}}$ point estimate ($+1.3\%$ ${\mathsf{WP}}$). But, the $90\%$ confidence interval of the estimated gain in win probability by attempting a field goal is $[-4\%,5\%]$, indicating that ${\mathsf{FG}}$ could either be a good or a bad decision. Also, about half of the bootstrapped models say ${\mathsf{Go}}$ is better. In other words, we do not have enough data to be confident in the win probability point estimates, and we don’t know the optimal fourth-down decision at this game-state. Further, in the bottom right plot, notice how most of the colors are light. This indicates that the optimal decision is uncertain at most other combinations of yardline and yards to go at this game-state.
@@
Example play 4. In Figure 15 we visualize the decision procedure for a fourth-down play in which the Ravens had the ball against the Jets in Week 1 of 2022. ${\mathsf{Punt}}$ provides a slight edge over ${\mathsf{Go}}$ according to the ${\mathsf{WP}}$ point estimate ($+0.08\%$ ${\mathsf{WP}}$). But most of the bootstrapped models say ${\mathsf{Punt}}$ is better and the $90\%$ confidence interval of the estimated gain in win probability by punting is $[-1\%,3\%]$, which is largely positive. So, even if the edge is small, we are confident in this edge and recommend that the Ravens should ${\mathsf{Punt}}$. Further, in the bottom right plot, notice how most of the colors are dark outside of a large white boundary region. This indicates that we have higher certainty in the estimated optimal decision at this game-state.
@@
Example play 5. In Figure 16 we visualize the decision procedure for an infamous fourth-down play in which the Raiders had the ball against the Rams in Week 14 of 2022. ${\mathsf{Go}}$ provides a strong edge over ${\mathsf{Punt}}$ according to the ${\mathsf{WP}}$ point estimate ($+4.1\%$ ${\mathsf{WP}}$). Further, many of the bootstrapped models say ${\mathsf{Go}}$ is better and the $90\%$ confidence interval of the estimated gain in win probability by going for it is $[-1\%,5\%]$, which is largely positive. Thus, we recommend that the Raiders should ${\mathsf{Go}}$666 In real life, the Raiders punted. Then, Rams quarterback Baker Mayfield countered with a successful 98-yard drive to win the game.

### v6

@@ Appendix C Tuning the number of bootstrapped datasets
We use ${\mathsf{boot\%}}$ to measure uncertainty in the estimated optimal decision, but we cannot necessarily rely on a non-technical football coach to fully process how ${\mathsf{boot\%}}$ maps to the strength of a recommendation. To facilitate easier communication to a coach, we bin decision recommendations into three buckets: confident (${\mathsf{boot\%}}\in[83\%,100\%]$), lean (${\mathsf{boot\%}}\in[67\%,83\%)$), and uncertain (${\mathsf{boot\%}}\in[0\%,67\%)$). Since we ultimately bucket decision recommendations into these three bins (see Section 4.2), we want the number of bootstrapped models $B$ to be large enough to stably categorize decisions into one of these three bins. We also want $B$ to be as small as possible in order to quickly evaluate a fourth-down recommendation during a football game. Hence, we conduct a stability analysis to choose $B$.
@@
$\begin{cases}p_{i}^{\text{confident, }(B)}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left({\mathsf{boot\%}}_{im}^{(B)}\in[83\%,100\%]\right),\\ p_{i}^{\text{lean, }(B)}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left({\mathsf{boot\%}}_{im}^{(B)}\in[67\%,83\%)\right),\\ p_{i}^{\text{uncertain, }(B)}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left({\mathsf{boot\%}}_{im}^{(B)}\in[0\%,67\%)\right),\\ p_{i}^{(B)}=\max\{p_{i}^{\text{confident, }(B)},\ p_{i}^{\text{lean, }(B)},\ p_{i}^{\text{uncertain, }(B)}\}.\end{cases}$ (29)
@@ Appendix D Baseline coaches’ decision model
To compare our decision-making procedure to the decisions that actual football coaches tend to make, we model the probability that a coach chooses a decision in $\{{\mathsf{Go}},{\mathsf{FG}},{\mathsf{Punt}}\}$ as a function of game-state. We use ${\mathsf{XGBoost}}$ to fit these coach probabilities. ${\mathsf{XGBoost}}$ works well here because we have $94,786$ fourth-down plays in our full dataset of plays since 1999, and each play is a reasonably approximate independent observation of a coach’s decision. In particular, we fit these coach probabilities as a function of yards to opponent endzone, yards to go, game seconds remaining, score differential, point spread, and era (1999-2001, 2002-2005, 2006-2013, 2014-2017, and 2018-present). In Figure 13 we visualize these coach decision models, and the results make intuitive sense. For the most part, coaches punt deep in their own territory and kick field goals near the opponent’s endzone, except for with one and sometimes two yards to go. Also, at the end of the game, coaches’ decision-making changes depending on the number of points they need to score to win the game.
Figure 13: Visualizing our model of the typical coach’s fourth-down decision as a function of yards to opponent endzone and yards to go, for various values of time remaining and score differential. Green, yellow, and red indicate that ${\mathsf{Go}}$, ${\mathsf{FG}}$, and ${\mathsf{Punt}}$ is the most likely decision, respectively. The color intensity reflects the likelihood that a coach makes that decision.
@@ Appendix E Additional example plays
Example play 3. In Figure 15 we visualize the decision procedure for a fourth-down play in which the Bears had the ball against the Jets in Week 12 of 2022. ${\mathsf{FG}}$ provides an edge over ${\mathsf{Go}}$ according to the ${\mathsf{WP}}$ point estimate ($+4.2\%$ ${\mathsf{WP}}$). But, the $90\%$ confidence interval of the estimated gain in win probability by attempting a field goal is $[-2\%,5\%]$, indicating that ${\mathsf{FG}}$ could either be a good or a bad decision. Also, $40\%$ of the bootstrapped models say ${\mathsf{Go}}$ is better. In other words, we do not have enough data to be confident in the win probability point estimates, and we don’t know the optimal fourth-down decision at this game-state. Further, in the bottom right plot, notice how most of the colors are light. This indicates that the optimal decision is uncertain at most other combinations of yardline and yards to go at this game-state.
@@
Example play 4. In Figure 16 we visualize the decision procedure for a fourth-down play in which the Ravens had the ball against the Jets in Week 1 of 2022. ${\mathsf{Punt}}$ provides a slight edge over ${\mathsf{Go}}$ according to the ${\mathsf{WP}}$ point estimate ($+1.3\%$ ${\mathsf{WP}}$). But most of the bootstrapped models say ${\mathsf{Punt}}$ is better and the $90\%$ confidence interval of the estimated gain in win probability by punting is $[0\%,3\%]$, which is positive. So, even if the edge is small, we are confident in this edge and recommend that the Ravens should ${\mathsf{Punt}}$. Further, in the bottom right plot, notice how most of the colors are dark outside of a large white boundary region. This indicates that we have higher certainty in the estimated optimal decision at this game-state.
@@
Example play 5. In Figure 17 we visualize the decision procedure for an infamous fourth-down play in which the Raiders had the ball against the Rams in Week 14 of 2022. ${\mathsf{Go}}$ provides a strong edge over ${\mathsf{Punt}}$ according to the ${\mathsf{WP}}$ point estimate ($+2.9\%$ ${\mathsf{WP}}$). Further, many of the bootstrapped models say ${\mathsf{Go}}$ is better and the $90\%$ confidence interval of the estimated gain in win probability by going for it is $[0\%,5\%]$, which is positive. Thus, we recommend that the Raiders should ${\mathsf{Go}}$999 In real life, the Raiders punted. Then, Rams quarterback Baker Mayfield countered with a successful 98-yard drive to win the game.